1.find an expression for 1/2||u+v||²+1/2||u-v||² in terms of ||u||²+||v||².

2.find an expression for ||u+v||²-||u-v||² in terms of u×v.

3.use the result of 2 to deduce an expression for ||u+v||² whenever u and v are orthogonal to each other.


1
Expert's answer
2021-06-18T10:53:57-0400

1. Given arbitrary vectors "u" and "v." We get


"||u+v||^2=(u+v)(u+v)"

"=(u,u)+(u,v)+(v,u)+(v,v)"

"=||u||^2+2(u,v)+||v||^2"



"||u-v||^2=(u-v)(u-v)"

"=(u,u)-(u,v)-(v,u)+(v,v)"

"=||u||^2+2(u,v)+||v||^2"

Then


"||u+v||^2+||u-v||^2"

"=||u||^2+2(u,v)+||v||^2+||u||^2-2(u,v)+||v||^2"

"=2||u||^2+2||v||^2"

Therefore


"\dfrac{1}{2}||u+v||^2+\dfrac{1}{2}||u-v||^2=||u||^2+||v||^2"

2.



"||u+v||^2-||u-v||^2"

"=||u||^2+2(u,v)+||v||^2-||u||^2+2(u,v)-||v||^2"

"=4(u,v)"

Therefore


"||u+v||^2-||u-v||^2=4u\cdot v"

3. If  u and v are orthogonal to each other, then "u\cdot v=0."

Hence


"||u+v||^2-||u-v||^2=0"

"=>||u+v||^2=|u-v||^2"




"\dfrac{1}{2}||u+v||^2+\dfrac{1}{2}||u-v||^2"

"=\dfrac{1}{2}||u+v||^2+\dfrac{1}{2}||u+v||^2"

"=||u+v||^2"

"=\dfrac{1}{2}||u-v||^2+\dfrac{1}{2}||u-v||^2"

"=||u-v||^2"


"=||u||^2+||v||^2"

"||u-v||^2=||u+v||^2=||u||^2+||v||^2"